OpenStudy (anonymous):

Find all real numbers x such that \[4x + 2 \ge 14\] and \[-21x + 1 > 22\] I am not looking for someone to give me the answer here, I am looking for someone to check my work to make sure I'm doing it right as I want to catch any mistakes now before I submit the assignment for grading. Posting the answers here will not get a medal!

OpenStudy (anonymous):

\[4x + 2 \ge 14\] is equal to \[x \ge 4\]

OpenStudy (anonymous):

-21x + 1 > 22 -1 -1 -21x > 21 /-21 /-21 x < -1

OpenStudy (anonymous):

How do you get x greater than or equal to 4?

OpenStudy (ajprincess):

What u have done for the second part is right:) Pls chck the first part

OpenStudy (anonymous):

There's a mistake in second part too

OpenStudy (ajprincess):

oops sorry my bad/. yup u r @Abdul_shabeer

OpenStudy (anonymous):

basically when I redid my work for the first part I got... \[x \ge 3\]

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

Okay and looking at the work for the 2nd one, I'm not really sure where I went wrong.

OpenStudy (anonymous):

Everything is fine except the final line

OpenStudy (anonymous):

21 / -21 = -1 correct?

OpenStudy (anonymous):

Okay, so when I was taught this, I was told that if your graphing these equations on a number line and you have to divide by a negative, the inequality sign flips to the opposite sign. Is that wrong?

OpenStudy (anonymous):

@abdul_shabeer If your graphing these equations on a number line and you have to divide by a negative, the inequality sign flips to the opposite sign. Is that wrong?

OpenStudy (anonymous):

If x is less than -1, then it doesn't satisfy the inequality

OpenStudy (anonymous):

So there would be no real numbers in this set?

OpenStudy (anonymous):

Wait, nevermind. We can still add the numbers from \[x \ge 3\] So it would be an open dot on 3 with the shading to the right?

OpenStudy (anonymous):

@ajprincess did I do it right?

OpenStudy (ajprincess):

Well I thnk u r right. cos real numbers include all positive negative integers , decimal numbers and all

OpenStudy (ajprincess):

\(x\ge3\) the dot also will be a shaded one

OpenStudy (ajprincess):

I mean ur second part too is right